OpenStudy (anonymous):

Factor the Following Polynomials. show work x³+64 10x²-6xy+35x-21y

OpenStudy (anonymous):

Use this formula: \[a^3 + b^3 = (a +b)(a^2 + b^2 - ab)\]

OpenStudy (anonymous):

You can write your expression as : \((x)^3 + (4)^3\).. Now use the above formula..

OpenStudy (anonymous):

someone walk me through this

OpenStudy (anonymous):

i would plug it in but idk which is a b act.

OpenStudy (anonymous):

ect.*

OpenStudy (anonymous):

I have written the formula above use that.. a =x and b = 4

OpenStudy (anonymous):

x^3+4^3=(x-4)(x^2 + 4^2 -x4)

OpenStudy (anonymous):

now distribute x - 4 to the parenthesis..

OpenStudy (anonymous):

what

OpenStudy (anonymous):

I said multiply (x - 4) with (x^2 + 16 - 4x)..

OpenStudy (anonymous):

ok so

OpenStudy (anonymous):

Or you can simply write it as such only: \[(x + 4)(x^2 + 16 - 4x)\] This will be your answer for first..

OpenStudy (anonymous):

\[x ^{2} * x-4= x^{3}\]

OpenStudy (anonymous):

+4

OpenStudy (anonymous):

or -4 idk

OpenStudy (anonymous):

No do not multiply it you are said to factorize so : \((x + 4)(x^2 + 16 - 4x)\) will be your answer for first...

OpenStudy (anonymous):

okay